Can someone please show me how to work out this problem? I need to show all work for this question which is to find the product.
(x^2+7) (x^3-5x-4)

Answer by ewatrrr(24785)

 
Hi, In 3 steps...
(x^2+7) (x^3-5x-4) = x^2(x^3-5x-4) + 7(x^3-5x-4)
= x^5-5x^3-4x^2 + 7x^3-35x-28
= x^5+2x^3-4x^2 -35x-28